Pleasee help me solve this equation:
Find the slope-intercept form of the equation of the line that passes through the point (6,5) and is parallel to the line 4x-3y=4.

Answer by mananth(16946)

4 x + -3 y = 4
Find the slope of this line
make y the subject
-3 y = -4 x + 4
Divide by -3
y = 1.33 x -1.33
Compare this equation with y=mx+b
slope m = 1.33 0.8
The slope of a line parallel to the above line will be the same
The slope of the required line will be 4/3
m= 4/3 ,point ( 6 , 5 )
Find b by plugging the values of m & the point in
y=mx+b
5 = 8.00 + b
b= -3
m= 4/3
Plug value of the slope and b in y = mx +b
The required equation is y = 4/3 x -3

m.ananth@hotmail.ca